OpenStudy (anonymous):

Plese help with this math problem. What is the value of n? -10=n/-0.2

OpenStudy (anonymous):

@zepdrix

OpenStudy (astrophysics):

\[-10 = \frac{ n }{ -0.2 }\]

OpenStudy (anonymous):

can u help?

OpenStudy (astrophysics):

Multiply both sides by -0.2

OpenStudy (astrophysics):

What happens/

HanAkoSolo (jamierox4ev3r):

sir astrophysics, you are one fast fellow. I commend you ^^ Carry on!

OpenStudy (astrophysics):

haha xD

OpenStudy (anonymous):

i got 2 by miltiplying -10 by-0.2

OpenStudy (astrophysics):

Yes exactly, but what happens to the other side?

OpenStudy (anonymous):

we dont know what n is so miltiply -0.2 by -0.2?

HanAkoSolo (jamierox4ev3r):



OpenStudy (astrophysics):

\[\color{red}{-.2} \times -10 = \frac{ n }{ -0.2 } \times \color{red}{-.2}\] notice how on the right side it gets cancelled out right?

OpenStudy (anonymous):

oh so the anserw is 20?

OpenStudy (astrophysics):

No it's 2, you were right

OpenStudy (anonymous):

so n=2

HanAkoSolo (jamierox4ev3r):

here's an example: \(\large\frac{x}{9} *9 =x\)

HanAkoSolo (jamierox4ev3r):

think of reciprocals :)

OpenStudy (astrophysics):

Yes,